the demand function for a product is D(q)=-2q2+60. use a difference with h=0.001 to estimate the marginal demand when 5 units are produced.

(A) $119.96 per unit

(B)$1 per unit

(C) -$0.04 per unit

(D) -$20 per unit

(E) -$40 per unit


1
Expert's answer
2022-03-14T13:25:28-0400

"\text{Marginal Demand(MD)}=\dfrac{dD}{dQ}\\
\text{MD}=-4q\\
MD_5=-4*5=-20"

Answer: (D) -$20 per unit